# Practical privacy metrics for synthetic data

### v1

@@ 4 Identifying disclosure from 1-way and 2-way relationships
As mentioned in the Introduction, what we can learn about disclosiveness of attributes can depend on our prior knowledge of the data set or the population from which it is drawn. It would not be practical to specify our prior probability for every possible combination of keys. However, checking two aspects of disclosure results can help us to check when we might have predicted the correct attribution with high probability without knowledge of all the keys. The first is a check for a target where a high proportion of records have one level of the target. The second is when there is a strong relationship between a target and one of the keys, so that one $tq$ pair accounts for many of the disclosive records. These two aspects are flagged by the values of check_1way and check_2way that are returned as part of the disclosure objects. Returning to the example in Section 2, we now use the function disclosure to get details for the two targets that were flagged as requiring checking; see Figure 1.
@@
R> print(d1_workab, to.print = c("check_1way"))
@@
   Level  All PctLevelAll totalDisclosive nLevelDis PctLevelDis
NO    NO 5000       88.64            2605      2482       95.28
We can see that it was the category "NO" of workab that contributed most to the disclosure risk; most survey respondents (89%) had never worked abroad. This level of the target accounted for 95% of apparent disclosures for workab. To predict this level for all of a group with the same $q$ could hardly be considered disclosive. This would be true if the intruder had access to the marginal distribution of workab, and even if they did not, some knowledge of the respondents’ background might suffice. The $tq$ pairs identified for workab both included the "NO" level of the target.
@@
R> print(d1_marital, to.print = c("check_2way"))
@@
This is a list with one for each of 4 syntheses
  target_key_levs npairs key key_target_total key_total PctTargetKeyLevel
7       SINGLE|19     10 age               91        92             98.91
4      MARRIED|41      5 age               59        73             80.82
6       SINGLE|18      5 age               91        92             98.91
8       SINGLE|22      5 age               85        91             93.41
For the target marital four $tq$ pairs are identified as contributing to large denominators. Those aged 19 are almost all SINGLE, and those aged 41 are mainly MARRIED. Again these would not be considered disclosive for an intruder with some knowledge of the respondents.
The thresholds for identifying one-way and two-way relationships can be modified. In each case there are two criteria that need to be satisfied, one number and one %. For one-way disclosure the parameter thresh_1way has the default value of $c(50,90)$, meaning that there must be at least 50 disclosive records for one level of a target and that $q$ values including this target must account for over 90% of all disclosive records. For two way relationships the thresh_2way has default value $c(4,80)$ and the algorithm first identifies all disclosive $tq$ combinations with denominators over 4. It then identifies the level of the key in $q$ with that best predicts this level of $t$ and checks if over 80% of the disclosive records with this key would have the correct prediction.
@@ 5 Excluding records
Having identified records where the apparent disclosure is something that would generally be known, one option is to exclude these key-target combinations explicitly from the measures. The following parameters for disclosure.summary and disclosure can be used for this.
@@
  • usetargetsNA(disclosure.summary) or usetargetNA (disclosure) ): Similar to the above for target(s) - note can be a vector for disclosure.summary.
  • exclude.keys, exclude.keylevs and exclude.targetlev: Three vectors of length the number of key-target pairs to be excluded from all tables 101010for disclosure.summary these parameters are supplied as lists with elements for each target in targets.
Figure 2: Plot from disclosure.summary for variables flagged by checks.
To illustrate exclusions we will use the Adult data set from the UCI machine learning repository Newman and Merz (1998) with almost 50 thousand records from the US Census income study, available as part of the arules package for R Hahsler et al. (2005). This is one of the data sets used in Giomi et al. (2022) to evaluate privacy risks for SD.Note that the two variables, texttcapital gain and texttcapital gain, originally numeric data, have been truncated so that they have only 123 and 99 distinct values repsectively. They have not been grouped further in the results presented here. Figure 2 is the output of disclosure.summary from the keys age, occupation, race and sex for the other 10 variables that are not in the keys.
Figure 2 is the output of disclosure.summary from the keys age, occupation, race and sex for the other 10 variables in the data. The disclosiveness of the GT data is relatievly low, compared to the previous example, because of the large sample size and the absence of any geographic identifiers. Three of the 10 variables capital.gain, capital.loss and native.country are flagged to check one-way relationships. The check1 column of the attribute table tells us that the levels contributing to disclosre are zeros for the first two and United-States for native.country. These levels make up 95%, 92% and 89% of all records. Three other variables are flagged as having disclosive two_way relationships workclass, marital, relationship with totals of 7,7 and 2 $tq$ combinations respectively. The summary function disclosure.summary does not allow target and key specific pairs to be excluded. To investigate this it is necessary to examine the output of disclosure111111see code in Appendix 3 for how to do this.. This showed that the largest contribution to check_2way for workclass was due to this being missing when occupation was missing, although other relationships between these two variables also contributed.
@@
Table 1 gives the results of excluding different entries in the tables of $q$ and $t$ from the attribute disclosure measures. Excluding the levels of the target flagged by textttcheck_1way reduces the $DiSCO$ to almost zero for these 3 variables. Adding exclusion of missing values reduced the disclosure for workclass and for some other variables a little. Adding the restriction to denominators of 1, reduced the disclosure for variables identified by check_2way to low levels. In the final columns we can see that restricting to denominators of 1, by itself, gives low levels of disclosure for all variables, with the exception of those flagged by check_1way. This approach was one that was trialed in an earlier version of these functions but discarded for giving misleading results for some data. It does not count some attribute disclosures that might well be found by an intruder. For example, if a small number of records with the same keys all had the same level of the target that corresponded to the level in the GT data. We feel that a better approach is to exclude specific target/key combinations.
@@ 6 Conclusions
The privacy metrics we propose here are in some sense the opposite of differential privacy (DP). DP claims to protect data from an intruder with arbitrary knowledge of the data, except of the one record that has the greatest influence on the likelihood of the results. In contrast our metrics require the user to specify keys that identify variables in the data that would expect to be known about individuals, as well as to specify the details of what they would expect an intruder to know about the data. Also, the routines flag cases where part of the disclosure measures come from one- or two-way relationships so that disclosure would be expected even in the absence of data.
We have evaluated the routines on a few data sets and set levels of the thresholds for this at what seem to be reasonable levels, but more experience with other data sets would be valuable.
We hope that these tools will be helpful to data holders who need to make decisions about the risks of releasing SD to the public, or to a restricted audience. They should also enable the disclosure risks of different synthesis methods to be evaluated.
@@ 7 Acknowledgement
Research Data Scotland (https://www.researchdata.scot/ funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package121212 from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control. We also thank the Scottish Centre for Administrative Data Research for continue to support the development of the synthpop package since its creation was supported by the ESRC funded SYLLS poject in 2012-14.
@@
Before defining the measures of identity and disclosure risk we need to introduce the notation that will be used to calculate them. The first step is to create the quasi-identifiers from the keys for the GT and SD. For the keys used in the example given in Section 2 the quasi-identifier that we will designate as $q$ for the first record in the GT data is:
@@
and that for the first record in the SD:
@@
In order to calculate identity disclosure measures, we need to compare the tables of $q$ from the GT and SD. For attribute disclosure measures we need to cross-tabulate $q$ with each target variable $t$ and compare findings from the SD with what would have been found from the GT data. In general, the levels of $q$ and sometimes $t$ in the GT and SD will not be the same. Before creating any tables, we need to define sets of $q$ and $t$ values that give the union of both sets of levels and align the tables so that their indices correspond.
For the GT data $d_{.q}$ is the count of records with the keys corresponding to the levels of $q$ and $q_{tq}$ is the count of records with this $q$ and level $t=1,...T$ of the target. The equivalent counts from the synthesised data are designated by $s_{.q}$ and $s_{tq}$. When a member of $q$ is in the GT data but not in the synthetic, $s_{.q}$ and $s_{tq}$ are all zero. Similarly when a member of $q$ is in the SD but not in the GT, $d_{.q}$ and $d_{tq}$ are all zero. The two tables can be written as shown in Table 1, where the total records in the GT data is $N_{d}$, made up of $N_{d~{}only}$ and $N_{d~{}both}$. The equivalent totals for the SD are $N_{s}$, $N_{s~{}only}$ and $N_{s~{}both}$.
@@
Table 2: Notation for tables from quasi-identifier ($q$) and target ($t$) from GT (upper table) and SD (lower table).
To calculate the % of records in the GT and SD we need:
@@
The intruder has information about the keys for an individual in the real data that they attempt to identify in the SD. They first attempt to find them in the SD, abd the % found is:
@@
Some of these records would not be unique in the SD, restricting to such records gives:
@@
To find an attribute from a set of keys, it is necessary to examine the distribution of $s_{tq}$ for groups defined by $q$. We define column proportions for the GT and SD as $pd_{tq}=d_{tq}/d_{.q}$ and for the synthetic as $ps_{tq}=s_{tq}/s_{.q}$.
Returning to the scenario described in Section 3.1, we must first define a measure of attribute disclosure for the GT data. This is based on the concept of l-diversity Machanavajjhala et al. (2006) that requires that each set of records defined by $q$ has at least $l(\geq{2}$ distinct values of the target. A data set is l2-diverse for $q$ and $t$ if all records for every $q$ have the same level of $t$ [^13] This could be generalised to $l>2$, but in practice the levels of targets are not generally exchangeable and a more practical approach would be to aggregate levels for certain targets. An attribute disclosure measure for the GT data can be defined as % Disclosive in Original :
@@
The equivalent measure for the SD, taken as if it were real, becomes :
@@
An intruder with access only to the SD, but with knowledge of $q$ from one or more individuals in the GT, would look them up in the SD. Some of their $q$ levels be key combinations that do not appear in the SD leaving the proportion that do appear as $iSO$ (in in Synthetic Original)
@@
A level of $q$ from an original record may identify more than one target in the SD, or identify the wrong target. To exclude these we require the records to be Disclosive in the SD and Correct when checked with the original giving:
@@
As we comment above the intruder would not be able to tell if records were identified as $DiSDiO$ rather than $DiSCO$, so we prefer the latter measure. However, the intruder can identify when the apparently disclosive record is not unique in the SD. This restriction can be imposed by requiring that the denominator for disclosive records in the SD does not exceed a 1, as described and discussed in Section 5.
@@
and the equivalent measure for the SD treated as if it were the original is
$CAPs=\sum\limits_{tq}{(ps_{tq}}s_{tq})/N_{s}.$
The measure DCAP is the percentage of $t$ correctly predicted in the SD $q$. This gives
@@
The $DCAP$ score can be expressed by scaling it from $baseCAPD$ to 1 (see Little et al. (2022); Pater and Smid (2024)), although this can result in some negative values. The $TCAP$ measure as defined by Little et al. (2022) uses the same count of disclosive records as $DISCO$, but uses as denominator the number of records that have keys represented in both the synthetic and the original data. It can also be scaled from $baseCAPD$ to 1, and as Little et al. (2022) demonstrate this measure can also take negative values.

### v2

@@ 4 Identifying disclosure from 1-way and 2-way relationships
As mentioned in the Introduction, what we can learn about disclosiveness of attributes can depend on our prior knowledge of the data set or the population from which it is drawn. It would not be practical to specify our prior probability for every possible combination of keys. However, checking two aspects of disclosure results can help us to check when we might have predicted the correct attribution with high probability without knowledge of all the keys. The first is a check for a target where a high proportion of records have one level of the target. The second is when there is a strong relationship between the target and one of the keys, so that one $tq$ pair accounts for many of the disclosive records. These two aspects are flagged by the values of check_1way and check_2way that are returned as part of objects of Returning to the example in 2 we now use the function disclosure to get details for the two targets that were flagged as requiring checking; see Figure 1.
@@
R> print(d1_workab, to.print = c("check_1way","check_2way"))
@@
  check Level totalDisclosive PctDisAll totalLevel pctdisLevel
1     1    NO            2661     53.22       2529       95.04
This is a list with one for each of syntheses
[[1]]
  total pctall max_denom    key        target_key pct_2way denom_2way
1     6    0.2         6    age           NO | 55     97.1        105
2     6    0.2         6 region NO | Podkarpackie     95.2        313
We can see that it was the category ”NO” of workab that contributed most to the disclosure risk; most survey respondents (89%) had never worked abroad. This level of the target accounted for 93% of apparent disclosures for workab. To predict this level for all of a group with the same $q$ could hardly be considered disclosive. This would be true if the intruder had access to the marginal distribution of workab, and even if they did not some knowledge of the respondents’ background might suffice. The $tq$ pairs identified for workab both included the ”NO” level of the target.
[…]
@@
R> print(d1_marital, to.print = c("check_1way","check_2way"))
@@
No 1 way checks for your target are flagged at your settings for thresh_1way
@@
This is a list with one for each of syntheses
[[1]]
  total pctall max_denom key   target_key pct_2way denom_2way
1     7    0.4         7 age MARRIED | 40     82.3         79
2     6    0.3         6 age  SINGLE | 20    100.0         79
For the target marital two $tq$ pairs are identified as contributing large denominators. Those aged 20 are all SINGLE, and those aged 40 are mainly MARRIED. Again these would not be considered disclosive for an intruder with some knowledge of the respondents.
The thresholds for identifying one-way and two-way relationships can be modified. In each case there are two criteria that need to be satisfied, one number and one %. For one-way disclosure the parameter thresh_1way has the default value of $c(50,90)$, meaning that there must be at least 50 disclosive records for one level of a target and that $q$ values including this target must account for over 90% of all disclosive records. For two way relationships the thresh_2way has default value $c(5,80)$ and the algorithm first identifies all disclosive $tq$ combinations with denominators over 5. It then identifies the level of the key in $q$ with that best predicts this level of $t$ and checks if over 80% of the disclosive records with this key would have the correct prediction.
@@ 5 Excluding records
Having identified records where the disclosure is not really telling us anything new, one option is to exclude some of these original records explicitly from the measures. The following parameters for multi.disclosure and disclosure can be used for this.
@@
  • usetargetsNA)(multi.disclosure) or usetargetNA (disclosure) ): Similar to the above for target(s) - note can be a vector for multi.disclosure.
  • exclude.keys, exclude.keylevs and exclude.targetlevFor disclosure only: Three vectors of length the number of key-target pairs to be excluded from all tables.
Figure 2: Plot from multi.disclosure for variables flagged by checks.
To illustrate exclusions we will use the Adult data set from the UCI machine learning repository [12] with almost 50 thousand records from the US Census income study, available as part of the arules package for R [8]. This is one of the data sets used in [6] to evaluate privacy risks for synthetic data. Figure 2 is the output of multi.disclosure from the keys age, occupation, race and sex for the other 10 variables in the data. The code used to calculate the results in this section is in Appendix 3. The disclosiveness of the original data is relatievly low, compared to the previous example, because of the large sample size and the absence of any geographic identifiers. Three of the 10 variables capital.gain, capital.loss and native.country are flagged to check one-way relationships. The check1 column of the attribute table tells us that the levels contributing to disclosre are zeros for the first two and United-States for native.country. These levels make up 95%, 92% and 89% of all records.
Three other variables are flagged as having disclosive two_way relationships workclass, marital, relationship with totals of 7,7 and 2 $tq$ combinations respectively. The summary function multi.disclosure does not allow target and key specific pairs to be excluded. To investigate this it is necessary to examine the output of disclosure999see code in Appendix 3 for how to do this.. This showed that the largest contribution to check_2way for workclass was due to this being missing when occupation was missing, although other relationships between these two variables also contributed.
@@
Table 1 gives the results of excluding different entries in the tables of $q$ and $t$ from the attrinute disclosure measures. Excluding the levels of the target flagged by textttcheck_1way reduces the $DiSCO$ to almost zero for these 3 variables. Adding exclusion of missing values reduced the disclosure for workclass and for some other variables a little. Adding the restriction to denominators of 1, reduced the disclosure for variables identified by check_2way to low levels. In the final columns we can see that restricting to denominators of 1, by itself, gives low levels of disclosure for all variables, with the exception of those flagged by check_1way. This approach was one that was trialed in an earlier version of these functions but discarded for giving misleading results for some data. It does not count some attribute disclosures that might well be found by an intruder. For example, if a small number of records with the same keys all had the same level of the target that corresponded to the level in the original data. We feel that a better approach is to exclude specific target/key combinations.
@@ 6 Conclusions
The privacy metrics we propose here are in some sense the opposite of differential privacy (DP). DP claims to protect data from an intruder with arbitrary knowledge of the data, except of the one record that has the greatest influence on the likelihood of the results. In contrast our metrics require the user to specify the details of what they would expect an intruder to know about the data. Most importantly, the user must specify keys that identify variables in the data that would expect to be known about individuals. Also, the routines flag cases where part of the disclosure measures come from one- or two-way relationships so that disclosure would be expected even in the absence of data. We have evaluated the routines on a few data sets and set levels of the thresholds for this at what seem to be reasonable levels, but more experience with other data sets would be valuable.
We hope that these tools will be helpful to data holders who need to make decisions about the risks of releasing synthetic data to the public, or to a restricted audience. They should also enable the disclosure risks of different synthesis methods to be evaluated.
@@ 7 Acknowledgement
Research Data Scotland (https://www.researchdata.scot/ funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package101010 from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control. We also thank the Scottish Centre for Administrative Data Research for continue to support the development of the synthpop package since its creation was supported by the ESRC funded SYLLS poject in 2012-14.
@@
Before defining the measures of identity and disclosure risk we need to introduce the notation that will be used to calculate them. The first step is to create the quasi-identifiers from the keys for the original and synthetic data. For the keys used in the example given in Section 2 the quasi-identifier that we will designate as $q$ for the first record in the original data is:
@@
and that for the first record in the synthetic data:
@@
In order to calculate identity disclosure measures, we need to compare the tables of $q$ from the original and synthetic data. For attribute disclosure measures we need to cross-tabulate $q$ with each target variable $t$ and compare findings from the synthetic data with what would have been found from the original data. In general, the levels of $q$ and sometimes $t$ in the original and synthetic data will not be the same. Before creating any tables, we need to define sets of $q$ and $t$ values that give the union of both sets of levels and align the tables so that their indices correspond.
For the original data $d_{.q}$ is the count of records with the keys corresponding to the levels of $q$ and $q_{tq}$ the count of records with this $q$ and level $t=1,...T$ of the target. The equivalent counts from the synthesised data are designated by $s_{.q}$ and $s_{tq}$. When a member of $q$ is in the original data but not in the synthetic, $s_{.q}$ and $s_{tq}$ are all zero. Similarly when a member of $q$ is in the synthetic data but not in the original, $d_{.q}$ and $d_{tq}$ are all zero. The two tables can be written as shown in Table 1, where the total records in the original data is $N_{d}$, made up of $N_{d~{}only}$ and $N_{d~{}both}$. The equivalent totals for the synthetic data are $N_{s}$, $N_{s~{}only}$ and $N_{s~{}both}$.
@@
Table 2: Notation for tables from quasi-identifier ($q$) and target ($t$) from original (upper table) and synthetic data (lower table).
To calculate the % of records in the original and synthetic data we need:
@@
The intruder has information about the keys for an individual in the real data that they attempt to identify in the synthetic data. They first attempt to find them in the synthetic data, abd the % found is:
@@
Some of these records would not be unique in the synthetic data, restricting to such records gives:
@@
To find an attribute from a set of keys, it is necessary to examine the distribution of $s_{tq}$ for groups defined by $q$. We define column proportions for the original and synthetic data as $pd_{tq}=d_{tq}/d_{.q}$ and for the synthetic as $ps_{tq}=s_{tq}/s_{.q}$.
Returning to the scenario described in Section 3.1, we must first define a measure of attribute disclosure for the original data. This is based on the concept of l-diversity [11] that requires that each set of records defined by $q$ has at least $l(\geq{2}$ distinct values of the target. A data set is l2-diverse for $q$ and $t$ if all records for every $q$ have the same level of $t$ [^11] This could be generalised to $l>2$, but in practice the levels of targets are not generally exchangeable and a more practical approach would be to aggregate levels for certain targets. An attribute disclosure measure for the original data can be defined as % Disclosive in Original :
@@
The equivalent measure for the synthetic data, taken as if it were real, becomes :
@@
An intruder with access only to the synthetic data, but with knowledge of $q$ from one or more individuals in the original, would look them up in the synthetic data. Some of their $q$ levels be key combinations that do not appear in the synthetic data leaving the proportion that do appear as $iSO$ (in in Synthetic Original)
@@
A level of $q$ from an original record may identify more than one target in the synthetic data, or identify the wrong target. To exclude these we require the records to be Disclosive in the Synthetic data and Correct when checked with the original giving:
@@
As we comment above the intruder would not be able to tell if records were identified as $DiSDiO$ rather than $DiSCO$, so we prefer the latter measure. However, the intruder can identify when the apparently disclosive record is not unique in the synthetic data. This restriction can be imposed by requiring the denominator in the synthetic data to exceed a 1, as described and discussed in Section 5.
@@
and the equivalent measure for the synthetic data treated as if it were the original is
$CAPd=\sum\limits_{tq}{(ps_{tq}}s_{tq})/N_{s}.$
The measure DCAP is the percentage of $t$ correctly predicted in the synthetic data $q$. This gives
@@
The $DCAP$ score can be expressed by scaling it from $baseCAPD$ to 1[10, 14], although this can result in some negative values. The $TCAP$ measure as defined by [10] uses the same count of disclosive records as $DISCO$, but uses as denominator the number of records that have keys represented in both the synthetic and the original data. It can also be scaled from $baseCAPD$ to 1, and as [10] demonstrate this measure can also take negative values.
@@
Appendix 3:: Code for analysis of the Adult data set.
library(arules)
data(AdultUCI)
names(AdultUCI) <- gsub("-",".", names(AdultUCI))
dim(AdultUCI)
##------ drop one of two identical education variables---------------------------
adultsample <- AdultUCI[,-c(5)]
names(adultsample)[-c(1,6,8,9)]
names(synadult\$syn)
##------ native.country moved to end as many levels that slow down synthesis-----
synadult <- syn(adultsample, visit.sequence = c(1:12,14,13) ,seed = 7868)
##------ get basic disclosure and make Figure 2----------------------------------
dd1 <- multi.disclosure(synadult,adultsample,
   keys = c("age" , "occupation","race", "sex" ))
pdf("vignettes/fig2dis.pdf",height=300, width=500)
print(dd1)
dev.off()
##------ exclude some target levels and missing values for targets---------------
dd2 <- multi.disclosure(synadult,adultsample,
    print.flag = FALSE, not.targetslev = c(rep("",5),"0","0","","United-States",""),
    keys = c("age" , "occupation","race", "sex" ))
##------ exclude some target levels and missing values for targets---------------
dd3 <- multi.disclosure(synadult,adultsample, print.flag = FALSE,
   not.targetslev = c(rep("",5),"0","0","","United-States",""),
    keys = c("age" , "occupation","race", "sex" ))
    dd1 <- multi.disclosure(synadult,adultsample, keys = c("age" , "occupation","race", "sex" ))
##-------restrict to denominators of 1 -----------------------------------------
dd4 <- multi.disclosure(synadult,adultsample, denom_lim =1,
    not.targetslev = c(rep("",5),"0","0","","United-States",""),
    usekeysNA = FALSE, usetargetsNA = FALSE,
    exclude_ov_denom_lim = TRUE,
    keys = c("age" , "occupation","race", "sex" ))
##----------just denominators not others---------------------------------------
dd5 <- multi.disclosure(synadult,adultsample, denom_lim =1,
    exclude_ov_denom_lim = TRUE,
    keys = c("age" , "occupation","race", "sex" ))
##-----------make table 1------------------------------------------------------
library(stringr)
cols12 <- dd1\$attrib.table[,1:2]
  dimnames(cols12)[[1]] <- word(dimnames(cols12)[[1]],2,2)
cols34 <- dd2\$attrib.table[,1:2]
  dimnames(cols34)[[1]] <- word(dimnames(cols34)[[1]],2,2)
cols56 <- dd3\$attrib.table[,1:2]
  dimnames(cols56)[[1]] <- word(dimnames(cols56)[[1]],2,2)
cols78 <- dd4\$attrib.table[,1:2]
  dimnames(cols78)[[1]] <- word(dimnames(cols78)[[1]],2,2)
cols910 <- dd5\$attrib.table[,1:2]
  dimnames(cols910)[[1]] <- word(dimnames(cols910)[[1]],2,2)
table.adult <- data.frame(cols12[order(dimnames(cols12)[[1]]),],
    cols34[order(dimnames(cols34)[[1]]),],
    cols56[order(dimnames(cols56)[[1]]),],
    cols78[order(dimnames(cols78)[[1]]),],
    cols910[order(dimnames(cols910)[[1]]),])
table.adult
##--------------------- check 2way for one target------------------------------
dis1 <- disclosure(synadult,adultsample, target = "workclass",
    keys = c("age" , "occupation","race", "sex" ))
print(dis1, to.print = c("attrib","check\_2way"))